OpenStudy (anonymous):

R is the first quadrant region enclosed by the x-axis, the curve y = 2x + b, and the line x = b, where b > 0. Find the value of b so that the area of the region R is 288 square units.

OpenStudy (anonymous):

the farthest i got was the integral from something to b of 2x+b

zepdrix (zepdrix):

When I first looked at this, I thought we would be integrating with these boundaries (the dots I made). But it appears they made things a little easier for us. Our region R is only in the first quadrant.

zepdrix (zepdrix):

So your "something" is simply x=0.

zepdrix (zepdrix):

\[\large\rm \int\limits_0^b 2x+b~dx=288\]Solve for b.

OpenStudy (anonymous):

and since b is a constant it goes out front..

OpenStudy (anonymous):

no wait thats wrong

zepdrix (zepdrix):

Ya, don't try to pull anything out front. Won't work out nicely :) Just power rule and stuff.

OpenStudy (anonymous):

oh god, well im getting b^2+b^2=288,

zepdrix (zepdrix):

Good good good.

OpenStudy (anonymous):

2b^2

OpenStudy (anonymous):

12

OpenStudy (anonymous):

i geuss i just didnt get that relation of x=0 tok fresh eyes to see it TY

zepdrix (zepdrix):

b=12, yay good job \c:/